OpenStudy (anonymous):

in a class of 43 students,the number of males is 8 less than twice the number of females.how many females and how many males are there in the class ?

OpenStudy (anonymous):

x = males, y = females x=2y-8 x+y=43 solve w/ augmented matrix / graphing / whatevs u want to x = 26, y = 17

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

m is males. m+8 = 2f so f is (m+8)/2 m+f = 43, get m and sub for y.

OpenStudy (anonymous):

Sub for f, sorry.

OpenStudy (anonymous):

number of female is 26 number f male is 17 male=2x-8 female=2x 2x+(2x-8)=43 2x+2x-8=43 4x-8=43 4x=51 x=12.8 so female is 2x=2(12.8) =26 male is 43-26=17